The Test score for a class are normally distributed . what is the the probability a student scored above 90. Given: µ = 75

Ơ = 10

1
Expert's answer
2022-05-17T00:12:53-0400

We have a normal distribution, "\mu=75, \sigma=10,x=90."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{90-75}{10}=1.5;
\\P(X>90)=P(Z>1.5)=1-P(Z<1.5)=\\
=1-0.9332=0.0668\text{ (from z-table)}."